OpenStudy (anonymous):

write the equation of a circle with its center at (5,-1)and radius 6

OpenStudy (anonymous):

Formula for a circle with centre (xo, yo) and radius r is: (x - xo)2 + (y - yo)2 = r2 Circle centre (0, 0) and radius 14: (x - 0)2 + (y - 0)2 = 142 x2 + y2 = 196

OpenStudy (anonymous):

A. (x+5)-(y-1)^2=6 B. (x-5)^2+(y+1)^2=36 C. (x-5)^2+(y-1)^2=36 D. (x+5)^2+(y-1)^2=6 @50_cent

OpenStudy (anonymous):

D

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

idk I do my homework

OpenStudy (anonymous):

I don't cheat

OpenStudy (anonymous):

give me best response

OpenStudy (anonymous):

what? all I wanna know is the work for it so that I can understand it

OpenStudy (anonymous):

go to youtube.com

OpenStudy (sirm3d):

let's start with the center radius equation of the circle as @50_cent gave \[(x-x_0)^2+(y-y_0)^2=r^2\] the center is \((x_0,y_0)\) and the radius is \(r\) what is the value of \(x_0,y_0, \text{ and } r\) the problem?

OpenStudy (anonymous):

x=5 y =1 r = 6. would it be C.?

OpenStudy (sirm3d):

almost. its x0 = 5, y0 = -1, and r = 6. now, sub these values into the equation, and you'll have the answer to your question.

OpenStudy (anonymous):

I got B. (x-5)^2+(y-1)^2=36

OpenStudy (sirm3d):

y-(-1) = y + 1 the correct answer is B.